OpenStudy (anonymous):

Solve: -2x - 13 = 8x + 7 help!!!

OpenStudy (bahrom7893):

-2x -8x = 13+7 -10x = 20 x = -2

OpenStudy (saifoo.khan):

-2x -8x = 7 + 13 -10 x = 20 x = -2

OpenStudy (anonymous):

saifoo, you seem to avoid all of the geometry questions for the easy ones. What's up with that?

OpenStudy (saifoo.khan):

lol, No i was kindda busy the whole night long with my website thing.. Thats y i didnt answered much. just came here to relax a lil.